# Safe and Economical UAV Trajectory Planning in Low-Altitude Airspace: A Hybrid DRL-LLM Algorithm with Compliance Awareness

### v2

@@ 7 Conclusion
In this paper, we have proposed a novel UAV trajectory planning algorithm tailored for low-altitude data collection in complex urban environments. Unlike prior works that focus on isolated objectives, our algorithm jointly addresses three key challenges: obstacle avoidance, regulatory compliance, and economically viable data collection. We have modeled the planning task as a partially observable Markov decision process and designed a hybrid optimization algorithm combining soft actor-critic algorithm with the large language model. Extensive experiments have demonstrated the effectiveness of our algorithm in improving data collection performance while reducing collisions, regulatory violations, and energy consumption.
Future work includes extending our algorithm to support multi-agent collaboration and real-world deployment in city-scale UAV networks. Moreover, integrating multi-modal sensor data and dynamically updating regulatory information will further enhance the generaliz ability and applicability of our algorithm.

### v3

@@ 7 Conclusion
This paper developed a hybrid unmanned aerial vehicle (UAV) trajectory planning algorithm that combines deep reinforcement learning with large language model reasoning for low-altitude data collection in complex urban environments. The proposed algorithm jointly considers obstacle avoidance, regulation awareness, and energy efficiency, achieving safe and adaptive decision-making under uncertainty.
Future work will explore extending the algorithm to multi-UAV collaboration and large-scale city scenarios. Another direction is to enhance three-dimensional obstacle avoidance and communication-aware coordination under partial or unstable network conditions. In addition, incorporating more realistic air-to-ground channel models, including probabilistic line-of-sight and non-line-of-sight propagation in urban environments, will help capture complex communication dynamics. Furthermore, integrating real sensor noise, multimodal perception, and dynamic regulatory feedback will be essential for improving robustness and real-world applicability in low-altitude economy networks.